a vessel containing an unknown volume of argon gas in cm3 at 300f and 101kpa was inverted and placed in cold ethanol . as the gas contracted the pressure was quadrulped while the temperature reduced by one-third . what is the final volume of the gas in dm3
1
Expert's answer
2019-05-01T04:28:35-0400

PV = nRT

T1 = 300 0F = (5/9 x (300 - 32)) + 273 = 422 K;

T2 = 300/3 = 100 0F = 311 K

P1 = 101 kPa = 101000 Pa

P2 = 101 kPa x 4 = 404 kPa = 404000 Pa

P1V1 / T1 = P2V2 / T2

101000 x V1 / 422 = 404000 x V2 / 311

V2 = 0.001 x V1